# A Survey on Graph Neural Network Acceleration: Algorithms, Systems, and Customized Hardware

### v1

@@ 6.3. Accelerators for General Workloads
HyGCN (Yan et al., 2020a) targets GCN, GraphSage, GINConv (Xu et al., 2018a), and DiffPool (Ying et al., 2018b) and treats their operations as two main stages, aggregation, and combination (Update in Eq. 1). Due to the distinctive execution patterns of the aggregation and combination stages, the proposed accelerator employs a separate processing engine for each stage and processes them in a dataflow manner. Specifically, the Aggregation engine utilizes parallel single-instruction multiple-data (SIMD) cores to exploit the intra-vertex parallelism resulting from the long feature vectors assigned to each vertex by the GNN algorithms. If the feature vector length is smaller than the total number of cores, the unassigned cores are allocated to other nodes, thereby enabling edge-level parallelism for this stage. As the graph connections are sparsely distributed, a sparsity eliminator is employed to identify a region in the adjacency matrix of the corresponding graph partition that stores the effective edges, using a window sliding and shrinkage technique. To further reduce the computation complexity, the aggregation operation is performed on a sampled set of neighbors. The sparsity eliminator and sampler also help to avoid fetching feature vectors of the nodes that are not connected to any edge in the current graph partition. Subsequently, the Combination engine treats the computation as a dense matrix-vector multiplication (MVM) and is realized as a group of systolic arrays that can work independently or cooperatively depending on the configuration. Each systolic array processes a small group of nodes, with the feature vectors (weights) of the nodes flowing through the row (column) dimension.
@@
BlockGNN (Zhou et al., 2021a) aims to compress the update stage (feature transformation) of GNN models for GraphSage-Pool (GS-Pool), GCN, GGCN (Marcheggiani and Titov, 2017), and GAT that require a large weight memory, high computing resources, and high computing latency when the feature size is large. To achieve this, they leverage block-circulant weight matrices, which are configured in the form of block-circulant matrices that can be accelerated by Fast Fourier Transform (FFT) and Inverse Fast Fourier Transform (IFFT). This transformation reduces the computation intensity and required storage for the update stage. The computation engine of BlockGNN is composed of two main units, the CirCore unit for the update stage and the VPU (vector processing unit) for the aggregation stage and non-linear operations. The CirCore unit is a three-stage pipeline, consisting of an FFT unit that converts data from the spatial domain to the spectral domain, a systolic array that performs element-wise product and accumulation, and an IFFT unit that transforms the results back to the spatial domain. The VPU is organized as a SIMD unit with m lanes, each with a parallel factor of 16.
@@
DeepBurning-GL (Liang et al., 2020a) is an automated framework designed to facilitate the development of hardware accelerators for GNNs. The framework takes a GNN model as input, which is designed using software frameworks such as DGL and PyG, and generates a hardware accelerator for the target FPGA platform. For the hardware design, the authors provide three templates. The first template is the GNN computation template that handles dense and sparse multiplications. Dense multiplications are implemented via a systolic array or a dot production array, while sparse multiplications are handled by an array of SIMD units or a ring-reduce topology (similar to EnGN). The second template is the memory template, which consists of a distributed on-chip buffer, cache, or degree-aware hierarchy on-chip buffer to support both regular and irregular memory accesses. The third template is the graph manipulation template, which supports graph sampling and reconstruction in various GNN models. To adapt to the specific characteristics of the target model and FPGA platform, the authors develop analytical models that analyze the processing requirements and adopt a design space exploration to tune hardware parameters for the required templates based on the graph properties, network architecture, and resource constraints. The authors validate their framework on various GNN algorithms, including GCN, GS-Pool, R-GCN, and EdgeConv (Wang et al., 2019a).
6.4. Accelerators for Special Workload
The literature in this field endeavors to direct their efforts toward a more specialized algorithm. More specifically, it concentrates on a singular GNN algorithm (here, mostly GCN), and subsequently tailors the proposed solutions accordingly. Given that the target model may contain multiple layers, the first subgroup develops a deep pipeline with custom layer modifications, which are elaborated upon in Section 6.4.1. Conversely, the second subgroup adopts a fixed hardware approach for all layers, which is described in detail in Section 6.4.2.
6.4.1. Layer-Customizable Deep Pipeline.
The studies presented in this section argue that a fundamental optimization in a GCN accelerator should involve reducing the expenses associated with memory transactions. To this end, they construct a deep pipeline that circumvents the need for off-chip (global) memory transactions concerning intermediate results among distinct GCN layers. Through the allocation of specific engines for each layer, they can additionally tailor hardware parameters based on the individual workload of each layer.
@@ 6.4. Accelerators for Special Workload
6.4.2. All-Layer Unified Architecture.
@@
Zhang et al. (Zhang et al., 2020b) focus on accelerating GCN inference for large graphs. The approach begins by partitioning the input data into smaller sizes that can fit into the on-chip resources for processing. Following this, a redundancy reduction scheme, such as GraphACT, and a node-reordering phase are applied to decrease global memory access. Upon completion of the pre-processing stage, the graph is passed to a hardware architecture that incorporates two primary modules: Aggregation and Weight Transformation. The Aggregation module employs parallel vector accumulators (VAs) for sparse matrix multiplication in the GCN aggregation process. Here, each VA parallelizes the feature dimension and different VAs process distinct edges. The Weight Transformation module leverages a systolic array to implement the update stage. These modules are bi-directionally connected to enable different modes of scheduling, where the order of executing the aggregation and update stage can vary based on the GCN and dataset properties.
@@
EnGN (Liang et al., 2020b) ASIC (TSMC 14nm with HBM 2.0 at 1GHz) 1802.9×1802.9$\times$1802.9 ×, 19.75×19.75$\times$19.75 ×, and 2.97×2.97$\times$2.97 × over CPU
cycle-accurate simulator (PyG &\&& DGL), GPU (PyG &\&& DGL), and HyGCN
Rubik  (Chen et al., 2021c) ASIC (45nm at 500MHz) 3.4246.7×3.42-46.7$\times$3.42 - 46.7 ×, and 1.314.16×1.3-14.16$\times$1.3 - 14.16 × over
HyGCN (Yan et al., 2020a) ASIC (TSMC 12nm with HBM at 1GHz) 1509×1509$\times$1509 ×, and 6.5×6.5$\times$6.5 × over
cycle-accurate simulator PyG-CPU, and PyG-GPU
FlowGNN (Sarkar et al., 2022) FPGA (Xilinx Alveo U50 at 300MHz) 24254×24-254$\times$24 - 254 ×, 1.3477×1.3-477$\times$1.3 - 477 ×, and 1.26×1.26$\times$1.26 × over
on-board evaluation PyG-CPU, PyG-GPU, and I-GCN
BlockGNN  (Zhou et al., 2021a) FPGA (Xilinx ZC706 at 100MHz) 2.3×2.3$\times$2.3 ×, and 4.28.3×4.2-8.3$\times$4.2 - 8.3 × over
DeepBurning -GL (Liang et al., 2020a) FPGA (Xilinx ZC706, KCU1500, and Alveo U50) 7.43346.98×7.43-346.98$\times$7.43 - 346.98 ×, 3.716.5×3.7-16.5$\times$3.7 - 16.5 ×, and 6.28×6.28$\times$6.28 × over
AWB-GCN (Geng et al., 2020) FPGA (Intel Stratix 10 SX) 3255×3255$\times$3255 ×, 80.3×80.3$\times$80.3 ×, 5.1×5.1$\times$5.1 × over
on-board evaluation PyG-CPU, PyG-GPU, and HyGCN
StreamGCN (Sohrabizadeh et al., 2022b) FPGA (Xilinx KU15P, Alveo U50 and U280) $8.2\times$ 18.2 ×, and 12.126.9×12.1-26.9$\times$12.1 - 26.9 × over
on-board evaluation at 201, 279, 290MHz PyG-CPU, and PyG-GPU
GraphACT (Zeng and Prasanna, 2020) FPGA (Xilinx Alveo U200 at 200MHz) (training speed) 1215×12-15$\times$12 - 15 ×, and 1.11.5×1.1-1.5$\times$1.1 - 1.5 × over
on-board evaluation CPU, and GPU
Zhang et al.  (Zhang et al., 2020b) FPGA (Xilinx Alveo U200 at 250MHz) 30×30$\times$30 ×, and 2×2$\times$2 × over
on-board evaluation CPU, and GPU - TensorFlow and C++ (Cuda)
BoostGCN (Zhang et al., 2021a) FPGA (Intel Stratix 10 GX 10M at 250MHz) 100×100$\times$100 ×, 30×30$\times$30 ×, $3$-45×45$\times$45 × over CPU (PyG &\&& DGL),
on-board evaluation GPU (PyG &\&& DGL), and FPGA ((Zhang et al., 2020b) &\&& HyGCN)
G-CoS  (Zhang et al., 2021b) FPGA (Xilinx VCU128 at 330 MHz) on-board evaluation 5.52×5.52$\times$5.52 ×, and 1.92×1.92$\times$1.92 × speedups over
HyGCN, AWB-GCN
GCoD  (You et al., 2022) FPGA (Xilinx VCU128 at 330 MHz) on-board evaluation 7.8×7.8$\times$7.8 ×, and 2.5×2.5$\times$2.5 × over HyGCN, AWB-GCN
I-GCN  (Geng et al., 2021) FPGA (Intel Stratix 10 SX at 330 MHz) on-board evaluation 368×368$\times$368 ×, 453×453$\times$453 ×, 16×16$\times$16 ×, and 5.7×5.7$\times$5.7 × over PyG-GPU, DGL-GPU, SIGMA, and AWB-GCN
Table 7 shows a review of accelerators, focusing on their validation methodology and reported speedup. These accelerators have been implemented on FPGAs or simulated as an ASIC design. A prevalent baseline for reporting the accelerator’s performance is CPU and GPU, with most works utilizing the state-of-the-art implementation of PyG for this purpose. The reported results may vary regarding whether they exclusively measure the runtime of the execution of the main kernel for their accelerator or include other associated overheads and present an end-to-end runtime.
7. GNN Acceleration: Special Graphs and Special GNNs
In the sections above we talked about acceleration techniques for the regular message-passing GNNs on regular graphs. However, many practical problems are better modeled with special graph types, e.g., heterogeneous graphs and dynamic graphs (both defined below), and solved with more sophisticated GNN architectures, e.g., heterogeneous GNNs and GNNs with attention. Many methods we discussed above may be applied to these special cases if their specialties are ignored, e.g, treating a heterogeneous graph as if it is homogeneous. However, the result would be suboptimal such as having high accuracy loss. In this section, we discuss techniques for handling these special cases especially.
7.1. GNN Acceleration on Heterogeneous Graphs
The graphs we talked about so far have nodes/edges of the same type, and they are called homogeneous graphs. The complex graphs in the real world are often associated with various types of nodes and edges (Dong et al., 2020; Javari et al., 2020; Huang et al., 2022), establishing a heterogeneous structure. For example in a citation graph, nodes can be authors, venues, and papers, and edges can denote an author published a paper, a paper published in a venue, etc. Different types of nodes and edges have substantially different meanings from each other and it is better to distinguish them explicitly. Heterogeneous graphs as a new data structure are defined below.
Definition 7.1 ().
A heterogeneous graph is defined as a directed graph ${\mathcal{G}}=({\mathcal{V}},{\mathcal{E}})$ associated with a node type mapping function $\phi:{\mathcal{V}}\rightarrow{\mathcal{A}}$ and an edge type mapping function $\tau:{\mathcal{E}}\rightarrow{\mathcal{R}}$. Each node $v\in{\mathcal{V}}$ belongs to one node type $\phi(v)\in{\mathcal{A}}$ and each edge $e\in{\mathcal{E}}$ belongs to one edge type $\tau(e)\in{\mathcal{R}}$.
Many GNNs are specially designed for heterogeneous graphs (Hu et al., 2020a; Zhang et al., 2019b; Long et al., 2021). They face similar scalability challenges as regular GNNs on homogeneous graphs and require acceleration. In the previous sections, we discussed acceleration methods that could be used for GNNs on heterogeneous graphs by disregarding their heterogeneity (i.e., different node and edge types). However, applying acceleration methods in a homogeneous manner by ignoring heterogeneity often leads to a significant decrease in performance. For example, for sampling algorithms, if a message passing step sample node neighbors uniformly without considering the node types, some infrequent types are likely missed out. Completely missing one type of node can downgrade the GNN performance a lot (Zhang et al., 2019b). A better sampling algorithm on heterogeneous graphs should perform a weighted sampling to include neighbor nodes of all types. Similarly for model pruning, heterogeneous GNNs often have separate model weights in the update function for each edge types (Schlichtkrull et al., 2018), and ignoring heterogeneity may prune the weights completely for some edge types, which makes the GNN unable to pass messages that type of edges and downgrade the performance. Therefore, accelerating GNNs while taking into account heterogeneity presents additional challenges and requires special treatment to capture heterogeneity. In the following, we will provide a summary of recent research on modeling and accelerating GNNs on heterogeneous graphs.
@@ 7.1. GNN Acceleration on Heterogeneous Graphs
HetGNN (Zhang et al., 2019b) is a GNN on heterogeneous graphs that introduces a novel heterogeneous neighbor sampler. The sampler belongs to the node-wise sampling category discussed in Section 3.2.1 but overcomes a special challenge on heterogeneous graphs. On heterogeneous graphs, node types in a target node’s receptive field can be very imbalanced, e.g., some types are very infrequent compared to other types. HetGNN argues that a regular sampler may miss some less common node types completely and result in insufficient node representations. Therefore, instead of randomly sampling, the HetGNN sampler first runs random walks with restart (RWR) in the neighborhood and then chooses nodes to form the computation graph according to their hit frequency by RWR. The sampler will also balance node types, such that the top frequent nodes of all types are chosen. With this RWR-based and type-balanced sampler, HetGNN can achieve fast training with good accuracy performance.
@@
HGT (Hu et al., 2020a) is a transformer-based GNN on heterogeneous graphs, which also introduces an efficient sampler to make the model handle Web-scale graph data with billions of edges. The HGT samples subgraphs in mini-batches. The sampler is able to 1) maintain a similar number of nodes and edges for each type in each subgraph and 2) keep the sampled subgraph dense. Specifically, for each node type it maintains a node budget and iteratively samples the same number of nodes with an importance sampling strategy to reduce variance. The sampling probability for each node type is calculated based on the square of the cumulative normalized degree of other nodes that have already been sampled in its budget. The model is shown to achieve state-of-the-art results on the Open Academic Graph (OAG) which contains millions of nodes and billions of edges.
@@
7.2. GNN Acceleration on Dynamic Graphs
Spatio-temporal data appears frequently in scientific research and real-world applications (Huang et al., 2020; Sanchez-Gonzalez et al., 2020). Dynamic graphs have shown great performance for modeling such spatio-temporal data to solve problems like molecular dynamics in biology (Yang et al., 2021a), information diffusion over social networks (Wang et al., 2021c), and epidemic spread in public health (Huang et al., 2021a). Dynamic graphs are graphs that can change over time. Changes include the addition and removal of nodes and edges to capture the evolving nature of the data. The evolving nature has imposed additional challenges for corresponding learning and reasoning on dynamic graphs as opposed to static graphs. Therefore, researchers have developed special spatio-temporal GNNs to jointly model the graph structural information on the spatial aspect and the evolving dynamics on the temporal aspect. The regular message passing in Equation (1) has been modified as is shown in Equation (18) for spatio-temporal GNNs, where the node representations are updated by both their spatial neighbors and their historical states.
Here ${\mathcal{E}}^{t-1}$ denotes the edges for graph at timestamp $t-1$, and sometimes the edges can be latent and need to be inferred first before the message passing.
Spatio-temporal GNNs are often applied to computation-heavy problems like the simulation of large-scale dynamical systems (Sanchez-Gonzalez et al., 2020), which have similar scalability concerns as regular GNNs on static graphs and require acceleration. However, spatio-temporal GNN acceleration incurs special challenges because model training needs to be done chronologically and graph operations (modification and sampling) need to consider both spatial and temporal neighbors. For example, the sampling methods introduced in Section 3.2 may be individually applied to dynamic graphs at each timestamp $t$, but they can fall short in performance as they only consider spatial properties of nodes but omit the temporal influence. In this section, we summarize recent works on spatio-temporal GNN acceleration, which are often applied for learning and reasoning over large-scale dynamic graphs.
TGL  (Zhou et al., 2022) proposes a unified framework for accelerating continuous-time temporal GNN training. It designs a temporal-CSR data structure for rapid access of temporal neighborhood candidates, and a parallel sampler is further equipped to accelerate temporal neighborhood sampling. Also, it proposes a novel random chunk scheduling technique for training acceleration where simply increasing the batch size would result in the obsolete node memory issue. TGL  (Zhou et al., 2022) shows its great capability in handling large-scale dynamic graphs where it achieves an average of 173$\times$ speedup on a multi-core CPU compared with the baselines.
@@ 7.2. GNN Acceleration on Dynamic Graphs
MTGNN (Wu et al., 2020a) applies GNNs to model the spatio-temporal graph for solving the multivariate time series prediction problem, where variables are dependent on each other but their dependence (graph structure) is latent. Naively assuming a fully-connected graph structure will result in $N^{2}$ edges and impose huge GNN computation time. MTGNN proposes a graph learning module to automatically extract the uni-directed relations among nodes, which greatly simplifies the structure and accelerates GNN computation. MTGNN also applies clustering to reduce memory usage by the intermediate states of nodes and save memory access time. Specifically, the edge search space is reduced by first randomly separating nodes into clusters and then only forming edges within each cluster. The clustering process for nodes is repeated in each iteration, which allows every pair of nodes in the same cluster to be possibly connected.
@@
GraphODEs (Huang et al., 2020, 2021a) combine the expressive power of GNNs with the principled ordinary differential equation (ODE) modeling for learning dynamical systems. LG-ODE (Huang et al., 2020) is a latent ODE generative model, which models multi-agent system dynamics with fixed graph structure in a continuous manner. The ODE function in LG-ODE is a GNN to capture the continuous interaction among agents, and the latent initial states for ODEs are inferred through another GNN-based encoder. Towards acceleration, existing discrete neural models learn a fixed-step transition function that takes the system state at time $t$ as input to predict the state at time $t+1$. An adaptive ODE solver will automatically adjust the times, where the graph interaction module is called to provide the balance between model performance and time consumption. Additionally, LG-ODE is especially superior at making long-range predictions. It is able to adjust its prediction length without re-feeding its previous predictions at the same fixed step during training, which greatly accelerates model inference. Moreover, since LG-ODE is only capable of handling dynamic graphs with evolving nodes but fixed edges, the follow-up work CG-ODE (Huang et al., 2021a) is proposed to jointly model the evolution of nodes and edges through two coupled ODE functions.
@@
Overall, spatio-temporal GNNs for dynamic graphs are extensions of static GNNs with special treatment for the temporal aspect. Traditional methods usually employ recurrent neural networks to capture the evolving nature of networks over time, but are usually time-consuming and fail to capture the long-term dependency. Recent advances to use positional encoding and neural ODEs, etc have shown promising results in speeding up dynamic GNNs. The works discussed above accelerate spatio-temporal GNNs in various ways to scale up them to computation-heavy dynamic system modeling and web-scale applications.
The proposed hardware accelerators in Section 6.2 have demonstrated the capability to support various graph topologies, provided that the graph is within the scale specified by the accelerator. This is attributed to the implementation of preprocessing techniques, such as graph re-ordering and tiling, within these works. However, challenges arise in supporting dynamic and heterogeneous graphs when the algorithm utilized to process them is altered. To address this issue without necessitating a complete redesign of the accelerator for each algorithmic change, software-hardware co-design can be extended. On the software front, an analyzer can be developed to evaluate the requirements of the graph and target algorithm and determine the optimal partitioning, mapping, and scheduling of operations to enable efficient execution on the hardware. On the hardware side, options include moving towards more general design architectures or supporting multiple specialized architectures that the software component can select based on the application requirements.
8. Future Directions
While significant progress has been achieved, the investigation into GNN acceleration remains in its early stages and has not kept pace with the rapid expansion of graph data. Numerous promising research avenues exist within this field, particularly in terms of acceleration from a COTS system and customized hardware perspective, which have only recently begun to garner attention. Additionally, there is a need to explore acceleration techniques tailored for heterogeneous and dynamic graphs. This section highlights several captivating research directions that warrant further exploration.
Better Graph Modification Algorithms Improving the efficiency of graph modification methods, either coarsening, sparsification, or condensation, represents a promising avenue for future research. In many graph datasets, nodes often have an excessive number of neighbors, which may provide redundant information or even noise. To accelerate GNN training, it may be possible to remove these neighbors more aggressively at the pre-processing stage. For example, through sparsification without preserving the global graph properties as practical GNNs are trained for preserving locality. Graph condensation is another promising technique that can enhance training efficiency, but it currently requires running the training process at least once to identify the best condensed graph. If the optimal graph can be determined earlier, e.g., identify a good condensed graph by training only a few epochs, this technique would become much more useful. Good graph modification methods can also benefit the automating designing and tuning GNN architectures. In practice, preprocessed simpler graphs will significantly accelerate the whole pipeline of GNN development.
Better Sampling Algorithms There are still several potential improvements for future research in sampling methods for GNNs. One potential direction is to investigate hierarchical sampling methods that can exploit the multi-scale structure of graphs to improve sampling efficiency and capture multi-scale features of graphs. Hierarchical sampling can be particularly useful for large-scale graphs with complex structures, such as social networks, where nodes can be organized into clusters or communities at different levels. Another possibility is to further investigate the sampling method for dynamic graphs, where nodes and edges change over time. Individually applying sampling to dynamic graphs at each timestamp is still not very efficient and fails to capture the temporal dynamics of nodes and edges on dynamic graphs.
Combined Inference Acceleration Algorithms Inference acceleration methods, such as pruning, quantization, and distillation, are often orthogonal to each other and can be combined to further improve the speed. Some initial steps have already been taken, such as leveraging distillation when ambitiously quantizing weights to binary. As the inference stage is more flexible training, e.g., no concerns like training stability and model convergence, these methods can be more easily combined in various ways. For instance, pruning and quantization can be combined to achieve a sparse model with low precision. Additionally, pruning can be used in conjunction with distillation to further refine the distilled student model.
9. Conclusion
In conclusion, we have presented a systematic view of acceleration methods for GNNs on large-scale datasets. Our survey covers GNN training and inference algorithms, efficient GNN systems, and customized hardware for GNNs. Additionally, we have discussed the special cases when GNNs are applied to heterogeneous graphs and dynamic graphs. Acceleration methods pave the way for more efficient GNN pipelines, enabling the use of larger GNN models for solving more complex tasks on graphs.
Acknowledgements.
We thank Linghao Song for the helpful discussions and comments throughout this work. We thank Xiang Song and Kezhao Huang for providing helpful feedback on an early draft of the paper. This work is partially supported by NSF (2211557, 1937599, 2119643), NASA, SRC, Okawa Foundation Grant, Amazon Research Awards, Cisco Research Grant, Picsart Gifts, and Snapchat Gifts.

### v2

@@ E.2. General Workloads: Dedicated Module per Stage
HyGCN (Yan et al., 2020b) targets GCN, GraphSage, GINConv (Xu et al., 2018a), and DiffPool (Ying et al., 2018b) and treats their operations as two main stages, aggregation, and combination (Update in Eq. 8). Due to the distinctive execution patterns of the aggregation and combination stages, the proposed accelerator employs a separate processing engine for each stage and processes them in a dataflow manner. Specifically, the Aggregation engine utilizes parallel single-instruction multiple-data (SIMD) cores to exploit the intra-vertex parallelism resulting from the long feature vectors assigned to each vertex by the GNN algorithms. If the feature vector length is smaller than the total number of cores, the unassigned cores are allocated to other nodes, thereby enabling edge-level parallelism for this stage. As the graph connections are sparsely distributed, a sparsity eliminator is employed to identify a region in the adjacency matrix of the corresponding graph partition that stores the effective edges, using a window sliding and shrinkage technique. To further reduce the computation complexity, the aggregation operation is performed on a sampled set of neighbors. The sparsity eliminator and sampler also help to avoid fetching feature vectors of the nodes that are not connected to any edge in the current graph partition. Subsequently, the Combination engine treats the computation as a dense matrix-vector multiplication (MVM) and is realized as a group of systolic arrays that can work independently or cooperatively depending on the configuration. Each systolic array processes a small group of nodes, with the feature vectors (weights) of the nodes flowing through the row (column) dimension.
@@
BlockGNN (Zhou et al., 2021c) aims to compress the update stage (feature transformation) of GNN models for GraphSage-Pool (GS-Pool), GCN, GGCN (Marcheggiani and Titov, 2017), and GAT that require a large weight memory, high computing resources, and high computing latency when the feature size is large. To achieve this, they leverage block-circulant weight matrices, which are configured in the form of block-circulant matrices that can be accelerated by Fast Fourier Transform (FFT) and Inverse Fast Fourier Transform (IFFT). This transformation reduces the computation intensity and required storage for the update stage. The computation engine of BlockGNN is composed of two main units, the CirCore unit for the update stage and the VPU (vector processing unit) for the aggregation stage and non-linear operations. The CirCore unit is a three-stage pipeline, consisting of an FFT unit that converts data from the spatial domain to the spectral domain, a systolic array that performs element-wise product and accumulation, and an IFFT unit that transforms the results back to the spatial domain. The VPU is organized as a SIMD unit with m lanes, each with a parallel factor of 16.
@@
DeepBurning-GL (Liang et al., 2020a) is an automated framework designed to facilitate the development of hardware accelerators for GNNs. The framework takes a GNN model as input, which is designed using software frameworks such as DGL and PyG, and generates a hardware accelerator for the target FPGA platform. For the hardware design, the authors provide three templates. The first template is the GNN computation template that handles dense and sparse multiplications. Dense multiplications are implemented via a systolic array or a dot production array, while sparse multiplications are handled by an array of SIMD units or a ring-reduce topology (similar to EnGN). The second template is the memory template, which consists of a distributed on-chip buffer, cache, or degree-aware hierarchy on-chip buffer to support both regular and irregular memory accesses. The third template is the graph manipulation template, which supports graph sampling and reconstruction in various GNN models. To adapt to the specific characteristics of the target model and FPGA platform, the authors develop analytical models that analyze the processing requirements and adopt a design space exploration to tune hardware parameters for the required templates based on the graph properties, network architecture, and resource constraints. The authors validate their framework on various GNN algorithms, including GCN, GS-Pool, R-GCN, and EdgeConv (Wang et al., 2019b).
E.2.1. Special Workload: Layer-Customizable Deep Pipeline.
@@
E.3. Special Workload: All-Layer Unified Architecture.
@@
Zhang et al. (Zhang et al., 2020) focus on accelerating GCN inference for large graphs. The approach begins by partitioning the input data into smaller sizes that can fit into the on-chip resources for processing. Following this, a redundancy reduction scheme, such as GraphACT, and a node-reordering phase are applied to decrease global memory access. Upon completion of the pre-processing stage, the graph is passed to a hardware architecture that incorporates two primary modules: Aggregation and Weight Transformation. The Aggregation module employs parallel vector accumulators (VAs) for sparse matrix multiplication in the GCN aggregation process. Here, each VA parallelizes the feature dimension and different VAs process distinct edges. The Weight Transformation module leverages a systolic array to implement the update stage. These modules are bi-directionally connected to enable different modes of scheduling, where the order of executing the aggregation and update stage can vary based on the GCN and dataset properties.
@@
Appendix F Detailed Discussion of Special Graphs and Special GNNs
F.1. Heterogeneous Graphs
@@
HetGNN (Zhang et al., 2019a) is a GNN on heterogeneous graphs that introduces a novel heterogeneous neighbor sampler. The sampler belongs to the node-wise sampling category discussed in Section 2.2.1 but overcomes a special challenge on heterogeneous graphs. On heterogeneous graphs, node types in a target node’s receptive field can be very imbalanced, e.g., some types are very infrequent compared to other types. HetGNN argues that a regular sampler may miss some less common node types completely and result in insufficient node representations. Therefore, instead of randomly sampling, the HetGNN sampler first runs random walks with restart (RWR) in the neighborhood and then chooses nodes to form the computation graph according to their hit frequency by RWR. The sampler will also balance node types, such that the top frequent nodes of all types are chosen. With this RWR-based and type-balanced sampler, HetGNN can achieve fast training with good accuracy performance.
@@
HGT (Hu et al., 2020c) is a transformer-based GNN on heterogeneous graphs, which also introduces an efficient sampler to make the model handle Web-scale graph data with billions of edges. The HGT samples subgraphs in mini-batches. The sampler is able to 1) maintain a similar number of nodes and edges for each type in each subgraph and 2) keep the sampled subgraph dense. Specifically, for each node type it maintains a node budget and iteratively samples the same number of nodes with an importance sampling strategy to reduce variance. The sampling probability for each node type is calculated based on the square of the cumulative normalized degree of other nodes that have already been sampled in its budget. The model is shown to achieve state-of-the-art results on the Open Academic Graph (OAG) which contains millions of nodes and billions of edges.
@@
F.2. Dynamic Graphs
[…]
@@ F.2. Dynamic Graphs
MTGNN (Wu et al., 2020b) applies GNNs to model the spatio-temporal graph for solving the multivariate time series prediction problem, where variables are dependent on each other but their dependence (graph structure) is latent. Naively assuming a fully-connected graph structure will result in $N^{2}$ edges and impose huge GNN computation time. MTGNN proposes a graph learning module to automatically extract the uni-directed relations among nodes, which greatly simplifies the structure and accelerates GNN computation. MTGNN also applies clustering to reduce memory usage by the intermediate states of nodes and save memory access time. Specifically, the edge search space is reduced by first randomly separating nodes into clusters and then only forming edges within each cluster. The clustering process for nodes is repeated in each iteration, which allows every pair of nodes in the same cluster to be possibly connected.
@@
GraphODEs (Huang et al., 2020, 2021b) combine the expressive power of GNNs with the principled ordinary differential equation (ODE) modeling for learning dynamical systems. LG-ODE (Huang et al., 2020) is a latent ODE generative model, which models multi-agent system dynamics with fixed graph structure in a continuous manner. The ODE function in LG-ODE is a GNN to capture the continuous interaction among agents, and the latent initial states for ODEs are inferred through another GNN-based encoder. Towards acceleration, existing discrete neural models learn a fixed-step transition function that takes the system state at time $t$ as input to predict the state at time $t+1$. An adaptive ODE solver will automatically adjust the times, where the graph interaction module is called to provide the balance between model performance and time consumption. Additionally, LG-ODE is especially superior at making long-range predictions. It is able to adjust its prediction length without re-feeding its previous predictions at the same fixed step during training, which greatly accelerates model inference. Moreover, since LG-ODE is only capable of handling dynamic graphs with evolving nodes but fixed edges, the follow-up work CG-ODE (Huang et al., 2021b) is proposed to jointly model the evolution of nodes and edges through two coupled ODE functions.